# 1]Univ. Artois, UR 4025, Laboratoire Systèmes Electrotechniques et Environnement (LSEE), F-62400 Béthune, France 2]Univ. Artois, UR 3926, Laboratoire de Génie Informatique et d’Automatique de l’Artois (LGI2A), F-62400 Béthune, France Time series classification with random convolution kernels based transforms: pooling operators and input representations matter

### v1

@@ 5.3 Sensitivity Analysis and Ablative Study
As previously shown, SelF-Rocket outperforms other existing methods based on random convolution kernels. In this section, the impact of the three key parameters on global accuracy performance is studied:
Following [21, 19], we choose 50 ‘development’ datasets from the 112 UCR datasets to conduct this analysis. We first remove the 12 datasets with the most training examples, then we select the sets that contain at least 5 examples per class and 100 training examples. From the remaining list, 50 were randomly selected. The complete list is available in Table 4.
The others parameters of SelF-Rocket, i.e. number of kernels, padding, dilation, bias, length of kernel, values inside the kernel, remain the same as the default parameters of MiniRocket.
@@ 5.3.1 Number of folds
The SelF-Rocket Feature Selection module employs a stratified -fold decomposition of the training set. We have conducted tests using the minimal number of folds possible, namely , up to a fixed value of . Figure 7 illustrates the impact of the number of folds on the classifier mean accuracy. No notable differences were observed when different values of were used.
@@ 5.3.2 Number of features
Figure 7 shows that the number of features has a relatively minor impact on the average accuracy of the final classifier.
5.3.3 Input Representations and Pooling Operators
| SelF-Rocket | SelF-Rocket | SelF-Rocket | SelF-Rocket | |
| ONLY BASE | ONLY DIFF | ONLY MIX | ||
| Win | – | 2 | 4 | 6 |
| Draw | – | 18 | 15 | 23 |
| Lose | – | 30 | 31 | 21 |
| Mean accuracy | 85.69 | 85.01 | 82.13 | 85.35 |
Table 2 provides a summary of the pairwise win/draw/loss between SelF-Rocket () and SelF-Rocket with only one set of representations.
The first-order difference is generally the least impactful representation, while the MIX set of representations demonstrates the best overall performance among the three.
| NO PPV | NO GMP | NO MIPV | NO MPV | NO LSPV | ||
| Win | – | 4 | 1 | 1 | 0 | 4 |
| Draw | – | 27 | 48 | 39 | 46 | 43 |
| Lose | – | 19 | 1 | 10 | 4 | 3 |
| Mean acc. | 85.69 | 85.44 | 85.63 | 85.36 | 85.60 | 85.67 |
The choice of a pooling operator or a set of pooling operators to generate features can affect the performance of the classifier in the classification process. Figure 9 and Table 3 show the impact on performance of not using one of the five pooling operators. It can be seen that not using GMP or LSPV has no significant impact on system performance. On the other hand, not using MIPV or PPV has a greater impact on overall performance.
@@ 5.4 Scalability and Compute Time Comparison
Compared to other ROCKET methods, SelF-Rocket incorporates a Feature Selection module between the global set of features and the final classifier. The time required for the generation of SelF-Rocket features is less than Hydra-MultiRocket, while the training of the classifier is similar in duration to that of MiniRocket.
Figure 10 shows the mean compute time for all resamples and all datasets of the Feature Selection module according to the number of folds and features. The classifier complexity depends on , the number of features and , the number of training examples. is a multiplier for the number of mini-classifiers, and impacts the complexity of a single classifier so the two parameters impact generally the compute time.
For each combination of IR and PO, we instantiate mini-classifiers to identify the optimal set of features. A RIDGE classifier with a complexity of is employed. In Figure 11(a) given that is a fixed value, namely 9,996, we can see that the mean time only depends on the number of training examples.
Figure 12(a) shows the average total computation time over the 50 development datasets, for all resamples. SelF-Rocket and MiniRocket training (Figure 12(a)) and prediction times (Figure 12(b)) are comparable. However, among the evaluated methods, MiniRocket remains the fastest overall. As illustrated in Figure 10, with and SelF-Rocket’s average feature selection time is around 1.3 seconds, which is significantly faster than its feature creation computation time. SelF-Rocket features take longer to create than MultiRocket features. This is due to the greater number of kernels that are instantiated. MultiRocket instantiates only 6,216 kernels for each of its input representations (DIFF and BASE). From these, features () are generated. In comparison, SelF-Rocket generates 9,996 kernels for each of these two representation types for a total of features ().
@@
HYDRA produces kernels : (number of kernels per group), (number of groups) and (maximum possible dilation value), with 2 features extracted per kernel (maximum and minimum responses). For each dataset, 512 ( and ) kernels per dilation are created with . In most of our cases will be between 7 and 10, so the total number of features generated will generally be between (7,128) and (10,240). This explains why HYDRA is the fastest in terms of classifier training and testing time. As illustrated in Figures 12(c) and 12(d), SelF-Rocket is faster than Hydra-MultiRocket in terms of both feature creation time and classifier testing time, with less discrepancy.
@@ 6 Conclusion and prospects
The two main contributions of this article are as follows:
First, we have shown that the choice of Input Representations and Pooling Operators significantly impacts the performance of TSC methods based on random convolution kernels. We have also shown that using multiple Pooling Operators, such as MultiRocket, leads to good performances on average, but is not optimal in specific cases: it is always preferable to select a single appropriate Pooling Operator.
Then, on this basis, we have proposed a new algorithm based on MiniRocket, SelF-Rocket, incorporating a feature selection stage. Experiments show that this algorithm selects the best Input Representations / Pooling Operator combination in most cases, leading to state of the art performances. In the tested implementation of SelF-Rocket, only 9,996 or 19,992 features are retained to train the RIDGE classifier, far fewer than competing methods such as MultiRocket or Hydra-MultiRocket, leading to fast predictions.
In this article, a minimal version of SelF-Rocket was evaluated, using 2 input representations, 5 pooling operators and a simple wrapper-based feature selection method. This implementation could be improved by:
Considering more input representations (second order difference, Fourier and Hilbert transforms, etc.),
Combining SelF-Rocket and HYDRA features,
Using HYDRA as a pooling operator within SelF-Rocket,
@@ Appendix A Additional Figures & Tables
| CricketX | CricketY | CricketZ |
| DistalPhalanxOutlineAgeGroup | DistalPhalanxOutlineCorrect | DistalPhalanxTW |
| Earthquakes | ECG200 | EOGHorizontalSignal |
| EOGVerticalSignal | EthanolLevel | FaceAll |
| Fish | FreezerRegularTrain | GunPointAgeSpan |
@@
| MiddlePhalanxOutlineCorrect | MixedShapesRegularTrain | MixedShapesSmallTrain |
| OSULeaf | Plane | PowerCons |
| ProximalPhalanxOutlineAgeGroup | ProximalPhalanxOutlineCorrect | ProximalPhalanxTW |
| ScreenType | SemgHandGenderCh2 | SemgHandMovementCh2 |
| SemgHandSubjectCh2 | ShapesAll | SmoothSubspace |
| SwedishLeaf | SyntheticControl | Trace |
| UWaveGestureLibraryX | UWaveGestureLibraryY | UWaveGestureLibraryZ |
| Worms | WormsTwoClass | |

### v4

@@ 5.3 Sensitivity Analysis and Ablative Study
In this section, we examine the influence of the five key parameters of the SelF-Rocket method on overall accuracy performance:
Following [16, 15], we choose 55 ‘development’ datasets from the 112 UCR datasets to conduct this analysis. We first select the set of datasets that contain at least 5 examples per class and 100 training examples. From that remaining list, 55 datasets were randomly selected. The complete list is available in Table 4.
5.3.1 Input Representations and Pooling Operators
| SelF-Rocket | SelF-Rocket | SelF-Rocket | SelF-Rocket | |
| ONLY BASE | ONLY DIFF | ONLY MIX | ||
| Win | – | 17 | 16 | 32 |
| Draw | – | 7 | 3 | 6 |
| Lose | – | 31 | 36 | 17 |
| Mean accuracy | 85.15 | 84.61 | 82.51 | 85.16 |
Table 2 provides a summary of the pairwise win/draw/loss between default SelF-Rocket (FULL) and SelF-Rocket with only one set of representations.
The first-order difference (DIFF) is generally the least impactful representation, while the MIX set of representations demonstrates the best overall performance among the three and is similar to the FULL version in terms of performance.
| NO PPV | NO GMP | NO MIPV | NO MPV | NO LSPV | ||
| Win | – | 19 | 7 | 15 | 17 | 21 |
| Draw | – | 6 | 46 | 29 | 16 | 24 |
| Lose | – | 30 | 2 | 11 | 22 | 10 |
| Mean acc. | 85.15 | 85.01 | 85.12 | 84.94 | 85.07 | 85.16 |
The choice of a pooling operator or a set of pooling operators to generate features can affect the performance of the classifier in the classification process. Figure 10 and Table 3 emphasize how not employing one of the five pooling operators affects performance. It can be seen that not using GMP or LSPV has no significant impact on performance, whereas not using PPV has a greater impact on overall performance.
@@ 5.3.2 Number of folds
[…]
@@ 5.3.3 Number of features
As exposed in Figure 12, the number of features has a relatively minor impact on the average accuracy of the final classifier.
5.3.4 Number of runs
The total number of decompositions of the initial data set depends on both the number of folds and the number of runs . The first one varies the distribution of the number of examples between the training set and the validation set, while the second one allows the original training set to be shuffled and then re-split differently. When , Figure 12 shows that the choice of the number of runs is relatively more important to improve the performance of the Features Selection Module.
5.3.5 Vote Validation
Once the highest median vote has been carried out, using the result of the vote directly can be risky, especially for small datasets because their mini-classifiers are trained on datasets that are too small when selecting the best set of features.
To avoid selecting a wrong combination that would generalize poorly, a vote validation system has been put in place by calculating the percentage of voters, who more or less agree with our final choice. If this percentage exceeds a certain value (threshold in Algorithm 2) then the vote is maintained, otherwise the vote is replaced by a default choice (PPV_MIX). Figure 13 displays the variation of the mean accuracy given a threshold value and a , e.g. if the , we calculate the percentage of voters with the final choice in their 4 best accuracy. In Figure 13, we can see that the higher the top, the higher the threshold thresh required to obtain greater accuracy. However, if the threshold is too high, the IR-PO combination selected is replaced too often by the default combination, which gives a lower accuracy.
A slight improvement in performance can be observed with the use of the vote validation system, but a greater impact of this system can be observed with smaller datasets cf. Appendix B.
@@ 5.4 Scalability and Compute Time Comparison
Compared to other ROCKET methods, SelF-Rocket includes a Feature Selection Module between the global feature set and the final classifier. The feature generation time of SelF-Rocket and Hydra + SelF-Rocket is similar to MultiRocket and Hydra + MultiRocket respectively, while the training time of SelF-Rocket’s Ridge classifier is similar to that of MiniRocket.
Figure 14 shows the mean computation time for all resamples and all datasets of the Feature Selection Module according to the number of runs and features. The classifier complexity depends on , the number of features and , the number of training examples. The number of runs is a multiplier for the number of mini-classifiers, and impacts the complexity of a single classifier so the two parameters impact generally the compute time. The comparison of the performance of SelF-Rocket (FULL) and SelF-Rocket (MIX) of Section 5.3.1 reveals that they are comparable. However, the feature selection phase of SelF-Rocket (MIX) is faster due to a lower number of IR-PO combinations (only 5). With , and , the SelF-Rocket Feature Selection Module takes 3.71 seconds on average for all resamples with the MIX version, compared to 11.12 seconds on average for the FULL version.
For each combination of IR and PO, we instantiate mini-classifiers to identify the optimal set of features. A Ridge classifier with a complexity of is employed. In Figure 15(a) given that is a fixed value, namely 5,000, we can see that the mean time only depends on the number of training examples. We have a value of (cf Algorithm 1), which explains why there is a cap on the Feature Selection Module’s time after this value.
In Figure 16(a), we can see the average total computation time over the 55 development datasets, for all resamples for the ROCKET methods. SelF-Rocket and MiniRocket Ridge classifier’s training (Figure 16(a)) and prediction times (Figure 16(b)) are comparable. However, among the evaluated methods, MiniRocket remains the fastest overall. SelF-Rocket features take longer to create than MultiRocket features. This is due to the greater number of kernels that are instantiated. MultiRocket instantiates only 6,216 kernels for each of its input representations (DIFF and BASE). From these, features () are generated. In comparison, SelF-Rocket generates 9,996 kernels for each of these two representation types for a total of features ().
@@
HYDRA produces kernels : (number of kernels per group), (number of groups) and (maximum possible dilation value), with 2 features extracted per kernel (maximum and minimum responses). For each dataset, 512 ( and ) kernels per dilation are created with . In most of our cases will be between 7 and 10, so the total number of features generated will generally be between (7,128) and (10,240). This explains why HYDRA is the fastest in terms of classifier training and testing time. As illustrated in Figures 16(c) and 16(d), Hydra + SelF-Rocket and SelF-Rocket are both faster than Hydra + MultiRocket and MultiRocket in terms of classifier testing time, with less discrepancy.
5.5 SelF-Rocket with an oracle
As shown in Section 5, selecting the best features can be difficult, particularly in small datasets. Therefore, the implementation of SelF-Rocket presented in this paper is not optimal. Knowing the optimal input representations and pooling operator, what would be the performances of SelF-Rocket?
Figure 17(a) illustrates the average mean accuracy ranking for the various classifiers previously tested, while Figure 17(b) presents the MCM of the Oracle version of SelF-Rocket and Hydra combined with SelF-Rocket, compared to HIVE-COTE v2.0 and other ROCKET methods. These results indicate the maximum potential performance of an algorithm based on SelF-Rocket. The question remains whether such performance is attainable, or if we can approach it with an improved Feature Selection Module.
| Mean-Accuracy | H + SelF-R (Oracle) 0.8916 | HC2 0.8914 | SelF-R (Oracle) 0.8892 | H + MR 0.8840 | MultiR 0.8814 | MiniR 0.8744 |
| H + SelF-R (Oracle) 0.8916 | Mean-Difference rc / r=c / rc Wilcoxon p-value | 0.0002 66 / 5 / 41 0.0236 | 0.0024 68 / 6 / 38 0.0008 | 0.0076 72 / 8 / 32 1e-04 | 0.0101 83 / 8 / 21 1e-04 | 0.0172 97 / 5 / 10 1e-04 |
| HC2 0.8914 | -0.0002 41 / 5 / 66 0.0236 | - | 0.0023 57 / 5 / 50 0.6579 | 0.0074 59 / 7 / 46 0.1046 | 0.0100 67 / 6 / 39 0.0051 | 0.0170 84 / 5 / 23 1e-04 |
| SelF-R (Oracle) 0.8892 | -0.0024 38 / 6 / 68 0.0008 | -0.0023 50 / 5 / 57 0.6579 | - | 0.0052 61 / 8 / 43 0.0261 | 0.0077 69 / 7 / 36 0.0001 | 0.0147 97 / 6 / 9 1e-04 |
| H + MR 0.8840 | -0.0076 32 / 8 / 72 1e-04 | -0.0074 46 / 7 / 59 0.1046 | -0.0052 43 / 8 / 61 0.0261 | - | 0.0026 68 / 9 / 35 0.0003 | 0.0096 80 / 5 / 27 1e-04 |
| MultiR 0.8814 | -0.0101 21 / 8 / 83 1e-04 | -0.0100 39 / 6 / 67 0.0051 | -0.0077 36 / 7 / 69 0.0001 | -0.0026 35 / 9 / 68 0.0003 | - | 0.0070 80 / 5 / 27 1e-04 |
| MiniR 0.8744 | -0.0172 10 / 5 / 97 1e-04 | -0.0170 23 / 5 / 84 1e-04 | -0.0147 9 / 6 / 97 1e-04 | -0.0096 27 / 5 / 80 1e-04 | -0.0070 27 / 5 / 80 1e-04 | If in bold, then p-value 0.05 |
@@ 6 Conclusion and prospects
This article presents a novel algorithm called SelF-Rocket, which builds upon MiniRocket by incorporating a feature selection stage.
The key idea behind SelF-Rocket is its dynamic selection of the best combination of input representations and pooling operators during the training phase. Experiments conducted on the UCR datasets show that SelF-Rocket, on average, selects a more effective combination of Input Representation and Pooling Operator compared to MiniRocket, and succeeds to achieve state-of-the-art performance. SelF-Rocket delivers performance comparable to MultiRocket (similarly, Hydra combined with SelF-Rocket matches the performance of Hydra combined with MultiRocket). However, SelF-Rocket offers the benefit of quicker classification prediction times due to its utilization of fewer features (9,996 or 19,992 for SelF-Rocket vs 50,000 for MultiRocket).
The version of SelF-Rocket evaluated in this study employs three sets of input representations, five pooling operators, and a simple wrapper-based feature selection method. Potential improvements to this implementation could include:
Incorporating additional input representations, such as second-order differences, Fourier transforms, and Hilbert transforms, among others.
Improving the vote validation systems could potentially bring performance closer to the optimal levels achieved by the oracle.
@@ Acknowledgments
The authors would like to thank Prof. Eamonn Keogh and all those who have contributed, and continue to contribute, to the maintenance of the University of California Riverside (UCR) TSC benchmark datasets and to open source implementations of the algorithms used in this article:
The original implementations of MiniRocket555https://github.com/angus924/minirocket, MultiRocket666https://github.com/ChangWeiTan/MultiRocket and Hydra777https://github.com/angus924/hydra were used as baseline.
Critical Difference diagrams were generated using aeon toolkit [23].
Multiple Comparison Matrices were generated using the original implementation888https://github.com/MSD-IRIMAS/Multi_Comparison_Matrix of [18].
The results provided by tsml-eval999https://github.com/time-series-machine-learning/tsml-eval were used to benchmark SelF-Rocket.
@@ Appendix A Additional Figures & Tables
| CricketX | CricketY | DistalPhalanxOutlineAgeGroup |
| DistalPhalanxOutlineCorrect | DistalPhalanxTW | Earthquakes |
| ECG200 | EOGHorizontalSignal | EOGVerticalSignal |
| EthanolLevel | FaceAll | Fish |
| FordA | FreezerRegularTrain | GunPointAgeSpan |
@@
| MiddlePhalanxOutlineCorrect | MiddlePhalanxTW | MixedShapesRegularTrain |
| MixedShapesSmallTrain | OSULeaf | PhalangesOutlinesCorrect |
| Plane | PowerCons | ProximalPhalanxOutlineAgeGroup |
| ProximalPhalanxOutlineCorrect | ProximalPhalanxTW | RefrigerationDevices |
| ScreenType | SemgHandMovementCh2 | SemgHandSubjectCh2 |
| ShapesAll | SmallKitchenAppliances | SmoothSubspace |
| StarLightCurves | Strawberry | SwedishLeaf |
| SyntheticControl | Trace | UWaveGestureLibraryX |
| UWaveGestureLibraryZ | Worms | WormsTwoClass |
| Yoga | ||
Appendix B Small vs Large datasets
We use all of the 36 UCR datasets with less than 100 training examples (thus missing from the sensitivity analysis) to see the impact of the number of runs 5.3.4 and vote validation 5.3.5 on small datasets. Figure 20 shows that these 2 SelF-Rocket components have an even greater effect on small datasets than on larger datasets. Regardless of the top considered, a significant consensus among voters is necessary to achieve optimal performance for the smallest datasets, as outlined in Figure 20(a). As Figure 20(b) illustrates it, using these two components together gets better results than using them separately.
The time needed to train SelF-Rocket depends on the training size. The larger the training size, the less important the time for SelF-Rocket feature selection becomes compared to the other time, as shown, for example, in Figure 21.